OpenStudy (mackenzie2013):

What is the slope of the line that passes through the points (-2, -3), and (5, 4)? -1 3 1 -3 Please Explain how to get answer!!

OpenStudy (anonymous):

The formula to get slope with two points is: y2-y1/x2-x1 4-(-3)=7 5-2=3 The slope is supposed to be 7/3, but i dont see it anywhere

OpenStudy (anonymous):

Did you put in the question correctly?

OpenStudy (mackenzie2013):

I copied and pasted it lol, Here it is again. I double checked! What is the slope of the line that passes through the points (-2, -3), and (5, 4)? -1 3 1 -3

OpenStudy (anonymous):

wait sorry, i made a mistake my bad.

OpenStudy (mackenzie2013):

Its alright.

OpenStudy (anonymous):

The formula to get slope with two points is: y2-y1/x2-x1 Work: 4-(-3)=7 5-(-2)=7 7/7=1 Answer: 1